Given the following information for XYZ Co., calculate the depreciation expense (in $): sales $50269; costs = $37937; addition to retained earnings = $2677; dividends paid = $1105; interest expense -$1401; tax rate - 29 percent.
Answer #1

Sales= $50, 269

all the expenses= $37,937

interest expense= $1401

Net revenue= 50,269- 37,937- 1401= $10,931

dividends paid + retained earnings= 1105+2677= $3782

the depreciation amount + tax amount = 10,931- 3782= $7149